OpenStudy (anonymous):

Write an equation in point-slope form for the line through the given point with the given slope. (8, 3); m = 6

OpenStudy (ajprincess):

(y-y1)/(x-x1)=m (y-3)/(x-8)=6 y-3=6(x-8) y-3=6x-48 y=6x-45